# ORBGRAND Is Exactly Capacity-achieving via Rank Companding

### v1

@@ III-B Proof of Main Result
$I(\mathsf{X};\mathsf{Y})$ $=\ln 2-\frac{1}{2}\int_{\mathcal{R}_{1}}\ln\Bigl(1+e^{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{+}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{2}}\ln\Bigl(1+e^{-\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{+}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{1}}\ln\Bigl(1+e^{-\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{-}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{2}}\ln\Bigl(1+e^{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{-}(y)\mathrm{d}y.$ (16)
@@ IV Application of CDF-ORBGRAND in BICM
High-order modulation has been a prevailing technique in contemporary communication systems, for achieving high spectral efficiency. In order to apply GRAND to high-order modulation, we study BICM systems, where a high-order coded modulation channel is effectively decomposed into multiple bit channels, with the decoding metric for each bit channel separately computed. Remarkably, we establish that CDF-ORBGRAND is capable of achieving the BICM capacity, which was initially derived in the literature by treating the decomposed bit channels as independent.
@@ IV-A BICM System Model
For notational convenience, $\underline{x}(w)$ is referred to as the codeword, and all subsequent descriptions are based on it. [^5] The original codeword $\tilde{\underline{x}}(w)$ can be recovered from $\underline{x}(w)$ through the de-interleaving operation. Since we consider random coding, elements of $\underline{x}(w)$ are i.i.d., uniformly drawn from $\{+1,-1\}$. After modulation, the $m$ bits $x_{i,1}(w),\ldots,x_{i,m}(w)$ as a block are mapped via the binary labeling function $\mu$ into $s_{i}(w)$, and the transmitted symbol sequence is denoted by
@@
$q_{j}^{+}(y)=\frac{1}{\lvert\mathcal{S}_{j}^{+}\rvert}\sum\limits_{s\in\mathcal{S}_{j}^{+}}p(y\lvert s),\quad q_{j}^{-}(y)=\frac{1}{\lvert\mathcal{S}_{j}^{-}\rvert}\sum\limits_{s\in\mathcal{S}_{j}^{-}}p(y\lvert s),$
@@
For $i=1,\cdots,N$ and $j=1,\cdots,m$, denote by $\mathsf{R}_{i,j}$ the rank of $\lvert\mathsf{T}_{i,j}\rvert$ among the sorted array consisting of $\left\{\lvert\mathsf{T}_{1,1}\rvert,\cdots,\lvert\mathsf{T}_{1,m}\rvert,\cdots,\lvert\mathsf{T}_{N,1}\rvert,\cdots,\lvert\mathsf{T}_{N,m}\rvert\right\}$, from $1$ (the smallest) to $mN$ (the largest). Let $\Psi_{1}(t),\cdots,\Psi_{m}(t)$ be the CDFs of $\lvert\mathsf{T}_{i,1}\rvert,\cdots,\lvert\mathsf{T}_{i,m}\rvert$ for $t\geq 0$, respectively. We further define the average CDF as
@@
We also introduce the sets $\mathcal{R}_{1,j}=\left\{y\lvert q_{j}^{+}(y)<q_{j}^{-}(y)\right\}$ and $\mathcal{R}_{2,j}=\left\{y\lvert q_{j}^{+}(y)>q_{j}^{-}(y)\right\}$.
@@ IV-B BICM Capacity
Analysis of BICM was initially conducted under the assumption of ideal interleaving [19], treating the BICM system model as $m$ mutually independent parallel binary-input memoryless channels, the $j$-th of which corresponds to the $j$-th bit of the labeling of the constellation symbols, with output pdf $q^{+}_{j}(y)$ under input $x=+1$ and $q^{-}_{j}(y)$ under input x=1x=-1. See Fig. 2 for an illustration.
Based on this parallel channel model, the BICM system achieves rate, i.e., the BICM capacity,
@@
Treating the $m$ bits within a symbol as independent results in a mismatched decoding rule, expressed as [26, 20]
@@ IV-C GRAND and Its Variants for BICM
for conducting the $q$-th query: if $P_{q,i,j}=-1$, the sign of $\kappa_{i,j}$ in $\underline{x}_{\text{hard}}$ is flipped, and otherwise it remains unchanged. If $\underline{x}_{\text{hard}}$ after flipping is a codeword, it is declared as the decoder output; otherwise, if all rows of $P$ are exhausted without finding a codeword, a decoding failure is declared.
For ORBGRAND, we arrange the rows of $P$ so that the sum reliability of the $q$-th row, $\sum_{i,j:P_{q,i,j}=-1}r_{i,j}$, is non-decreasing with $q$, where $r_{i,j}$ is the rank of $\lvert t_{i,j}\rvert$. Similar to Section II-B, we have the following unified decoding rule for GRAND algorithms:
@@ IV-D Rate Analysis
$I_{\text{ORB,BICM}}$ $=m\ln 2-\inf\limits_{\theta<0}\Bigg\{\sum_{j=1}^{m}\Big(\int_{0}^{1}\ln(1+e^{\theta t})\mathrm{d}t$
$-\frac{\theta}{2}\int_{\mathcal{R}_{1,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{+}(y)\mathrm{d}y$
$-\frac{\theta}{2}\int_{\mathcal{R}_{2,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{-}(y)\mathrm{d}y\Big)\Bigg\}$ (26)
@@ IV-E Numerical Results
The numerical results are shown in Fig. 4. It can be observed that ORBGRAND incurs a significant gap to the BICM capacity under both Gray and SP labelings. For example, with 8PSK and Gray labeling at SNR 15 dB, $I_{\text{ORB,BICM}}$ falls short of the BICM capacity (i.e., $I_{\text{CDF-ORB,BICM}}$) by approximately 0.5 bits per channel use. Moreover, this gap grows as the modulation order increases, reflecting the heightened sensitivity of ORBGRAND to high-order modulation in fading channels. Under the same conditions, the gap with SP labeling is even more pronounced than that with Gray labeling, indicating the impact of the choice of labeling scheme.
The performance gain of CDF-ORBGRAND over ORBGRAND can be attributed to the improved alignment of decoding metrics via rank companding, as commented at the beginning of Section III. This phenomenon appears to be more evident in the intermediate regime of SNR, and as modulation order increases.
In Fig. 4 we also include the achievable rates $\tilde{I}_{\text{ORB,BICM}}$ of BICM derived in [25] under the assumption of ideal interleaving; [^6] Similar to Fig. 1, here we rectify an error in the numerical program in [25] and update the numerical results. see the end of Appendix E for details. Although the derivation is approximate, we observe that the curves of $\tilde{I}_{\text{ORB,BICM}}$ are very close to the exact ones of ${I}_{\text{ORB,BICM}}$, indicating that the ideal interleaving assumption is usually valid in practice.
@@
Figure 4: Achievable rates of ORBGRAND, CDF-ORBGRAND, and BICM capacity for QPSK, 8PSK, and 16QAM over Rayleigh fading channel with perfect CSI. Line styles indicate curve type: dashed for ORBGRAND with ideal interleaving, dash-dot for ORBGRAND with non-ideal interleaving, solid for CDF-ORBGRAND and BICM capacity. Line colors indicate modulation format: black for QPSK, blue for 8PSK, red for 16QAM.
@@ V Conclusion
In this work, we first analyze the achievable rates of ORBGRAND and its variant under rank companding, CDF-ORBGRAND, from the perspective of the GMI for general binary-input memoryless channels. An somewhat unexpected yet fundamental finding is that the achievable rate of CDF-ORBGRAND is identical to the symmetric capacity, indicating that rank companding loses no information asymptotically. Building upon this finding, we extend the analysis to BICM and demonstrate that, CDF-ORBGRAND also achieves the BICM capacity. As numerical experiments reveal, ORBGRAND experiences a substantial rate gap in fading channels.
We emphasize that our analytical findings are information-theoretic in nature, which becomes effective as code lengths grow large. In practical scenarios with finite code lengths, CDF-ORBGRAND still exhibits a performance gap relative to ML decoding. Moreover, various practical optimization techniques for ORBGRAND have been proposed to enhance its performance in realistic settings; see, e.g., [7, 6, 10, 17, 18]. Exploring these methods and evaluating CDF-ORBGRAND under finite code lengths will be the focus of our future work.
@@ Appendix A Proofs of Lemmas 1, 4, and 7
$\quad\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\Big].$ (27)
@@
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\Big]$
$=\quad$ $\frac{1}{2}\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1\Big]$
$+\frac{1}{2}\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}>0\right)\Big\lvert\mathsf{X}_{i}(1)=-1\Big].$ (28)
@@
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1\Big]$
$=\quad$ $\mathbb{E}\Big[\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}\Big]\Big],$ (29)
@@
$\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}=y\Big]$
$=$ $\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\cdot\mathbf{1}(q^{+}(y)<q^{-}(y))\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}=y\Big]$
@@
Note that $\mathsf{R}_{i}$ in (30) is the rank of $|t|=\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert$ when inserted into a sorted array of N1N-1 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y)$ the upper branch of (30). Hence we have
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1\Big]$
@@
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\Big]$

### v2

@@ III-B Proof of Main Result
$I(\mathsf{X};\mathsf{Y})$ $=\ln 2-\frac{1}{2}\int_{\mathcal{R}_{1}}\ln\left(1+e^{\bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\bigr\rvert}\right)q^{+}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{2}}\ln\left(1+e^{-\bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\bigr\rvert}\right)q^{+}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{1}}\ln\left(1+e^{-\bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\bigr\rvert}\right)q^{-}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{2}}\ln\left(1+e^{\bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\bigr\rvert}\right)q^{-}(y)\mathrm{d}y.$ (16)
@@ IV Application of CDF-ORBGRAND in BICM
High-order modulation has been a prevailing technique in contemporary communication systems, for achieving high spectral efficiency. In order to apply GRAND to high-order modulation, we study BICM systems, where a high-order coded modulation is effectively decomposed into multiple bit channels, with the decoding metric for each bit channel separately computed. Remarkably, we establish that CDF-ORBGRAND is capable of achieving the BICM capacity, which was initially derived in the literature by treating the decomposed bit channels as independent.
@@ IV-A BICM System Model
For notational convenience, in the sequel, $\underline{x}(w)$ will be referred to as the codeword, and the original codeword $\tilde{\underline{x}}(w)$, which can be recovered from $\underline{x}(w)$ through a de-interleaving operation, will not be needed in the analysis. Since we consider random coding, elements of $\underline{x}(w)$ are i.i.d., uniformly drawn from $\{+1,-1\}$. After modulation, the $m$ bits $x_{i,1}(w),\ldots,x_{i,m}(w)$ as a block are mapped via the binary labeling function $\mu$ into a symbol $s_{i}(w)$, and the transmitted symbol sequence is denoted by
@@
$q_{j}^{+}(y)=\frac{1}{\lvert\mathcal{S}_{j}^{+}\rvert}\sum\limits_{s\in\mathcal{S}_{j}^{+}}p(y|s),\quad q_{j}^{-}(y)=\frac{1}{\lvert\mathcal{S}_{j}^{-}\rvert}\sum\limits_{s\in\mathcal{S}_{j}^{-}}p(y|s),$
@@
For $i=1,\cdots,N$ and $j=1,\cdots,m$, denote by $\mathsf{R}_{i,j}$ the rank of $\lvert\mathsf{T}_{i,j}\rvert$ among the sorted array consisting of $\left\{\lvert\mathsf{T}_{1,1}\rvert,\cdots,\lvert\mathsf{T}_{1,m}\rvert,\cdots,\lvert\mathsf{T}_{N,1}\rvert,\cdots,\lvert\mathsf{T}_{N,m}\rvert\right\}$, from $1$ (the smallest) to $mN$ (the largest). Let $\Psi_{1}(t),\cdots,\Psi_{m}(t)$ be the CDFs of $\lvert\mathsf{T}_{\cdot,1}\rvert,\cdots,\lvert\mathsf{T}_{\cdot,m}\rvert$ for $t\geq 0$, respectively. We further define the average CDF as
@@
We also introduce the sets $\mathcal{R}_{1,j}=\left\{y|q_{j}^{+}(y)<q_{j}^{-}(y)\right\}$ and $\mathcal{R}_{2,j}=\left\{y|q_{j}^{+}(y)>q_{j}^{-}(y)\right\}$.
@@ IV-B BICM Capacity
Analysis of BICM was initially conducted under the assumption of ideal interleaving [20], treating the BICM system model as $m$ mutually independent parallel binary-input memoryless channels, the $j$-th one corresponding to the $j$-th bit of the labeling of the constellation symbols, with output pdf $q^{+}_{j}(y)$ under input $x=+1$ and $q^{-}_{j}(y)$ under input x=1x=-1; see Fig. 2 for an illustration.
Based on this parallel channel model, the BICM system achieves the rate, i.e., the BICM capacity,
@@
Treating the $m$ bits within a symbol as independent is in fact a mismatched decoding rule, which can be expressed as [27, 21]
@@ IV-C GRAND and Its Variants for BICM
for conducting the $q$-th query: if $P_{q,i,j}=-1$, the sign of $\kappa_{i,j}$ in $\underline{x}_{\text{hard}}$ is flipped, and otherwise it remains unchanged. If $\underline{x}_{\text{hard}}$ after flipping is a codeword, it is declared as the decoder output. If all rows of $P$ are exhausted without finding a codeword, a decoding failure is declared. For ORBGRAND, we arrange the rows of $P$ so that the sum reliability of the $q$-th row, $\sum_{i,j:P_{q,i,j}=-1}r_{i,j}$, is non-decreasing with $q$, where $r_{i,j}$ is the rank of $\lvert t_{i,j}\rvert$.
Similar to Section II-B, we have the following unified decoding rule for GRAND algorithms:
@@ IV-D Rate Analysis
$I_{\text{ORB,BICM}}$ $=m\ln 2-\inf\limits_{\theta<0}\Bigg\{\sum_{j=1}^{m}\Biggl(\int_{0}^{1}\ln(1+e^{\theta t})\mathrm{d}t$
$-\frac{\theta}{2}\int_{\mathcal{R}_{1,j}}\bar{\Psi}\left(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\right)q_{j}^{+}(y)\mathrm{d}y$
$-\frac{\theta}{2}\int_{\mathcal{R}_{2,j}}\bar{\Psi}\left(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\right)q_{j}^{-}(y)\mathrm{d}y\Biggr)\Bigg\}$ (26)
@@ IV-E Numerical Results
The numerical results are shown in Fig. 4. It can be observed that ORBGRAND incurs a significant gap to the BICM capacity under both Gray and SP labelings. For example, with 8PSK and Gray labeling at SNR 15 dB, $I_{\text{ORB,BICM}}$ falls short of the BICM capacity (i.e., $I_{\text{CDF-ORB,BICM}}$) by approximately 0.5 bits per channel use. Moreover, this gap grows as the modulation order increases, reflecting the heightened sensitivity of ORBGRAND to high-order modulation in fading channels. Under the same conditions, the gap with SP labeling is slightly more pronounced than that with Gray labeling, indicating the impact of the choice of labeling scheme.
The performance gain of CDF-ORBGRAND over ORBGRAND can be attributed to the improved alignment of decoding metrics via rank companding, as commented at the beginning of Section III. This phenomenon appears to be the most evident in the intermediate regime of SNR, and as modulation order increases.
In Fig. 4 we also include the achievable rates $\tilde{I}_{\text{ORB,BICM}}$ of BICM derived in [25] which was derived under the assumption of ideal interleaving; [^5] Similar to Fig. 1, here we rectify an error in the numerical program in [25] and update the numerical results. see the end of Appendix E for details. Although the derivation relies upon an ideal assumption, we observe that the curves of $\tilde{I}_{\text{ORB,BICM}}$ are very close to the exact ones of ${I}_{\text{ORB,BICM}}$, indicating that the ideal interleaving assumption is usually sufficient for practical purposes.
@@
Figure 4: Achievable rates of ORBGRAND, CDF-ORBGRAND, and BICM capacity for QPSK, 8PSK, and 16QAM over Rayleigh fading channel with perfect CSI. Line styles indicate curve type: solid for CDF-ORBGRAND and BICM capacity, dash-dot for ORBGRAND (26), and dashed for ORBGRAND under the assumption of ideal interleaving (64) in [25]. Line colors indicate modulation format: black for QPSK, blue for 8PSK, red for 16QAM.
@@ V Conclusion
In this work, we first analyze the achievable rates of ORBGRAND and its variant under rank companding, CDF-ORBGRAND, from the perspective of the GMI for general binary-input memoryless channels. A somewhat unexpected yet basic finding is that the achievable rate of CDF-ORBGRAND is identical to the symmetric capacity, indicating that rank companding loses no information asymptotically. Building upon this finding, we extend the analysis to BICM and demonstrate that, CDF-ORBGRAND also achieves the BICM capacity. By contrast, ORBGRAND sometimes experiences a substantial rate loss in fading channels.
We emphasize that our analytical findings are information-theoretic in nature, which become effective as code lengths grow large. In practical scenarios with finite code lengths, CDF-ORBGRAND still exhibits a performance gap relative to ML decoding. Meanwhile, various practical optimization techniques for ORBGRAND have been proposed to enhance its performance in realistic settings; see, e.g., [8, 7, 11, 18, 19]. In consideration of these, evaluating CDF-ORBGRAND under finite code lengths will be the focus of our future work.
@@ Appendix A Proofs of Lemmas 1, 4, and 7
$\quad\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\right].$ (27)
@@
$\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\right]$
$=$ $\frac{1}{2}\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Bigg|\mathsf{X}_{i}(1)=+1\right]$
$+\frac{1}{2}\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}>0\right)\Bigg|\mathsf{X}_{i}(1)=-1\right].$ (28)
@@
$\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Bigg|\mathsf{X}_{i}(1)=+1\right]$
$=$ $\mathbb{E}\left[\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}\right]\right],$ (29)
@@
$\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}=y\right]$
$=$ $\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\cdot\mathbf{1}(q^{+}(y)<q^{-}(y))\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}=y\right]$
@@
Note that $\mathsf{R}_{i}$ in (30) is the rank of $\lvert t\rvert=\bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\bigr\rvert$ when inserted into a sorted array of N1N-1 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y)$ the upper branch of (30). Hence we have
$\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Bigg|\mathsf{X}_{i}(1)=+1\right]$
@@
$\mathbb{E}\left[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\right]$